Mackmyra Whisky AB needs additional funds for further expansion. The problem with their business of producing and selling whisky is a long cash conversion cycle, which means that cash comes in only 9, 12 or 15 years after production due to the long storage time of the whisky in oak barrels.

Mackmyra AB considers to take a loan of 30 million SEK for the expansion at a borrowing rate of 3%. However, Mackmyra's new CFO, a former JIBS student, proposes to sell 30-liter whisky barrels in advance to customers right after production. (Customers would take care themselves regarding the storage of their whisky barrels for some years until the whisky is bottlened and ready for consumption.) However, the price of a freshly distilled whisky barrel (called a zero-year barrel) is lower than those of 9, 12 or 15 years old barrels. The CFO estimates that Mackmyra could sell the zero-year barrel at a price of 17750 kr. Currently, 9-year barrels are sold at a price of 25000 kr, 12-year barrels are sold at a price of 28500 kr and 15-year barrels are sold at a price of 31000 kr. (For the sake of simplicity, we assume that there won't be price changes for the barrels in the coming years.)

a.) Calculate the NPV of this proposal. Assumption: From three produced whisky barrels, one is sold as 9-year, one is sold as 12-year, and the third one is sold as a 15-year barrel.

(Hint: assume that selling three zero-year barrels implies to loose the income from selling them as 9-, 12-, and 15-year barrel later on.)

b.) At which price of the zero-year barrel is Mackmyra indifferent between taking a loan or selling the zero-year barrels? (or to put it differently, at which price of the zero-year barrel do we have a break-even?)

c.) Consider now the perspective of an investor who is interested to buy zero-year barrels at a price of 17750 kr. What is the IRR of buying three zero-year barrels and keeping them until 9, 12 and 15 years respectively, and to sell them in those years at the prices stated above? (We assume again that prices do not change and that the investor has no storage costs.)

d) If the investor has an opportunity cost of capital of 3%, should she buy the zero-year barrel at a price of 17750 kr?
Answer #1

Part (a)

Let's assume Mackmyra sells 3 barrels of the zero-year barrel at a price of 17750 kr. He / she therefore loses the opportunity to sell at a future date,1 barrel each of 9-year barrels at a price of 25000 kr, 12-year barrels at a price of 28500 kr and 15-year barrels at a price of 31000 kr.

So, the cash flows will look like as shown below in the table. Please be guided by the second column in the table titled “Linkage” to understand the mathematics. That explains the mathematics behind each row. The last row highlighted in yellow is your answer. Figures in parenthesis mean negative values. All financials below are in KR.

Year, N Linkage 0 9 12 15
Sale of 3 barrels of Zero year A = 17750 x 3      53,250
Opportunity loss B = 1 x Price of each       (25,000)       (28,500)        (31,000)
Net cash flows C = A + B      53,250       (25,000)       (28,500)        (31,000)
Discount rate R 3%
Discount factor DF = (1 + R)-N      1.0000         0.7664         0.7014         0.6419
PV of cash flows PV = DF x C      53,250       (19,160)       (19,989)        (19,898)
NPV Sum of all PV      (5,797)

Hence, NPV of the proposal is - 5,797 Kr.

Part (b)

Let P be the break even price. From the table above, PV of future cash flows are = -19,160 - 19,989 - 19,898 (I have highlighted these values in the bold in the table above) = - 59,047.

Hence, we need to solve the equation: NPV = 3 x P - 59,047 = 0

Hence, P = 59,047 / 3 =  19,682.49 KR per barrel

Part (c)

The IRR calculation is shown below. I have produced the snapshot from my model. The last row highlighted in yellow is your answer. In the adjacent cell highlighted i blue, I have shown the excel formula used the yellow colored cell to get the output.

X Cash flows (53,250) 187 Year 188 0 189 1 10 190 2 0 191 0 192 4 0 193 194 6 195 7 10 196 8 0 25,000 197 198 10 199 11 0 28,

IRR = 3.89%

Part (d)

IRR = 3.89% > opportunity cost of capital of 3%, she should buy the zero-year barrel at a price of 17750 kr.